OpenStudy (samigupta8):

the no. of integral values of a for which x^2-(a-1)x+3=0 has both roots positive and x^2+3x+6-a=0 has both roots negative is a.0 b.1 c.2

OpenStudy (anonymous):

consider the roots as 'a' and 'b' then in first part a>0 ,b>0 so a+b>0 => a-1>0 a>1 put d>=0 think of same codn in 2nd part

OpenStudy (anonymous):

consider roots be r and t rather than a and b

OpenStudy (anonymous):

kya kehna h?

OpenStudy (samigupta8):

mtlabb..

OpenStudy (anonymous):

matlab samjhi ya nahi?

OpenStudy (samigupta8):

one equation is a^2-2a-13>=0

OpenStudy (anonymous):

take intersection of both the inequalities

OpenStudy (samigupta8):

second equation is 4a-15>=0

OpenStudy (ranga):

Let p and q be the roots of x^2-(a-1)x+3=0. For both roots to be positive (p+q) = (a-1) > 0; ==> a > 1 and pq = 3 > 0 which we already know it is. But there is one more condition to be met. The discriminant must be positive too. There are cases where p + q > 0 and pq > 0 and yet the roots might be complex. Discriminant = (a-1)^2 - 12 > 0 too. Find the intersection of those 2 inequalities for this equation.

OpenStudy (samigupta8):

how am i going to find intersection a>1 and a^2-2a-13>=0

OpenStudy (ranga):

(a-1)^2 - 12 > 0 (a-1)^2 > 12 (a-1) > sqrt(12) a > sqrt(12) + 1

OpenStudy (samigupta8):

intersection wud be a>sqrt 12+1

OpenStudy (ranga):

yes.

OpenStudy (ranga):

do the similar procedure for x^2+3x+6-a=0 has both roots negative

OpenStudy (samigupta8):

then wt abt second one

OpenStudy (samigupta8):

i got value of a as a belongs to (15/4,6)

OpenStudy (anonymous):

@ranga you'll get some other roots if you'll take sqrt by that way

OpenStudy (ranga):

x^2+3x+6-a=0 Assume r and s are the roots. r + s = -3 and rs = 6 - a If both roots are negative r + s < 0 which it is because r + s = -3 rs should be positive. So (6-a) > 0 or a < 6. Also the discriminant = 9 - 4(6-a) > 0. Find the intersection of these two inequalities as well and then intersect with the previous solution.

OpenStudy (samigupta8):

i m also geting dis part only

OpenStudy (ranga):

@divu.mkr (a-1)^2 > 12 ==> (a-1) > sqrt(12) or (a-1) < - sqrt(12). The second one can be thrown away because we already have one inequality where a > 1 and it will not intersect with (a-1) < - sqrt(12)

OpenStudy (anonymous):

ohhhkk

OpenStudy (ranga):

1 + sqrt(12) = 4.46. So 4.46 < a < 6 The only integer value for a is 5.

OpenStudy (ranga):

^^^ is this for a different problem or for the original posting?

OpenStudy (anonymous):

same

OpenStudy (ranga):

That looks like a solution for x^2 - 2x - 11 = 0

OpenStudy (anonymous):

yep...she was having a prob in solving that

OpenStudy (samigupta8):

thanku...for telling me the solution and bearing my madness for such a long tym..... truly thanks....dhanyavaad

OpenStudy (ranga):

I had to look up dhanyavaad :)